Fowler, Inc., just paid a dividend of $2.55 per share on its stock. The dividends are expected to grow at a constant rate of 3.9 percent per year, indefinitely. If investors require a return of 10.4 percent on this stock, what is the current price? What will the price be in three years? In 15 years?
Answer #1

Part A:

P0 = D1 / (Ke - g)

P0 = Current Price

D1 = Expected Div after 1 Year

Ke = COst of Equity

g = Growth Rate

D1 = D0(1+g)

= $ 2.55 (1+0.039)

= $ 2.55(1.039)

= $ 2.6495

P0 = D1 / (Ke - g)

= $ 2.6495 / (0.104-0.039)

= $ 2.6495 / 0.065

= $ 40.76

Part B:

P3 = D4 / (Ke - g)

P3 = Price after 3 Years

D4 = Expected Div after 4 Years

Ke = COst of Equity

g = Growth Rate

D4 = D0(1+g)^4

= $ 2.55 (1+0.039)^4

= $ 2.55(1.039)^4

= $2.55 * 1.1654

= $ 2.9717

P3 = D4 / (Ke - g)

= $ 2.9717 / (0.104-0.039)

= $ 2.9717 / 0.065

= $ 45.72

Part c:

P15 = D16 / (Ke - g)

P15 = Price after 15 Years

D16 = Expected Div after 16 Years

Ke = COst of Equity

g = Growth Rate

D16 = D0(1+g)^16

= $ 2.55 (1+0.039)^16

= $ 2.55(1.039)^16

= $2.55 * 1.8444

= $ 4.7031

P15 = D16 / (Ke - g)

= $ 4.7031 / (0.104-0.039)

= $ 4.7031 / 0.065

= $ 72.36